OpenStudy (anonymous):

What is the remainder when (3x^4 + 2x^3 – x^2 + 2x – 24) ÷ (x + 2) ?

OpenStudy (anonymous):

I used synthetic division and got -12

OpenStudy (anonymous):

I appreciate the help, but this is a multiple choice question and -12 isn't an option.

OpenStudy (anonymous):

dang sorry

OpenStudy (anonymous):

put x= -2 in (3x^4 + 2x^3 – x^2 + 2x – 24) and the answer will be 3*16+2* (-8) - 4 -4 -24 = 48-16-32 = 0 So, the remainder is zero